OpenStudy (anonymous):

Find the radius of a circle with the given circumference. 15pi ft.

OpenStudy (anonymous):

radius is 7.5 or 15/2 ft

OpenStudy (anonymous):

circumference= diameter * pi., so your diameter is 15. Radius is half the diameter, so your radius is 7.5 ft

OpenStudy (anonymous):

2*pi*r = 15*pi 2r = (15*pi)/pi 2r = 15 r = 15/2 r = 7.5 ft